what is the true equation for 2a+5b=3 is it (1/2, _4/5) or (0, 3/5). Can you show me how you came up with the answer
Answer by stanbon(75887)

what is the true equation for 2a+5b=3 is it (1/2, _4/5) or (0, 3/5). Can you show me how you came up with the answer
-----------
To test (1/2, -4/5) substitute 1/2 in place of a and -4/5 in place of b
and see if you get a true statement:
2(1/2) + 5(-4/5) = 3
1-4 = 3
-3=3
No, that is not true.
----------------------
To test (0,3/5) substitute 0 in place of a and 3/5 in place of b.
2*0 + 5*(3/5) = 3
0 + 3 = 3
3= 3
Yes, that is true.
So (0,3/5) is the solution for the equation.
=============
Cheers,
Stan H.